OpenStudy (anonymous):

taking the derivative of (1+1/x)^lnx

OpenStudy (anonymous):

\[(1+\frac{ 1 }{ x })^{\ln x}\]

OpenStudy (anonymous):

take the log, take the derivative, then multiply by this again

OpenStudy (anonymous):

\[(1+\frac{1}{x})\ln(\ln(x))\] is when you get when you take the log then take the derivative using product and chain rule, then multiply your result by \[(1+\frac{1}{x})^{\ln(x)}\]